OpenStudy (anonymous):

the nth term of a sequence o 4n+1 write down the first three terms of the sequence ??? i dont get it someone explain please

OpenStudy (anonymous):

Ok the first term is when n=1, second is when n=2, third when n=3 and so on... so term 1 is 4(1)+1=5, term 2 is 4(2)+1=9, term 3 is 4(3)+1=13...

OpenStudy (anonymous):

ooooh oooh oooh i get it thanks